# Generation of near GeV protons by tightly focused laser interacting with down-ramp density plasma

### v1

@@ Abstract
Enhancing proton energy is of great importance in laser-driven proton acceleration with finite laser energy for applications such as cancer therapy. We demonstrate an unusual acceleration scheme that achieves higher proton energies at lower laser energies by reducing the focal spot size. Through particle-in-cell simulations and theoretical modeling, we find that at small spot sizes (0.8 m), the proton energy is enhanced by 83.5%, much higher than that under conventional spot sizes (3 m). This is because the proton acceleration is dominated by electrons driven by an enhanced ponderomotive force at small spot sizes, generating stronger charge-separation fields that propagate faster. To further improve the proton energy, we analytically derive an optimal electron density profile, which enables phase-stable proton acceleration with an energy increased by 60%. These results are robust across parameter variations, suggesting that advanced focusing techniques and optimal plasma profiles could loose the requirement of laser energy, potentially reducing the dependence on large-scale laser facilities for medical and scientific applications.
@@ II Results
We perform a series of two-dimensional particle-in-cell (PIC) simulations using the EPOCH epoch code. A Gaussian laser pulse with a normalized amplitude of , a pulse duration of 30 fs and a wavelength of 800 nm is incident normally on a fully ionized hydrogen target with a density of 20 and a thickness of 7.5 m, where is the classical critical density. The simulation domain is a 40 m × 10 m rectangular box divided into 4000 × 1000 cells, with 20 macro-particles per cell per species. To investigate the effect of laser spot size, we vary the full width at half maximum (FWHM) of the laser focal spot from 0.8 m to 5 m, with representative cases of 0.8 m and 3 m selected for detailed analysis.
@@ II.1 Tightly focused laser
Figure 1(a) shows the temporal evolution of the maximum proton energy for the 0.8 m and 3 m spot sizes, and the proton energy spectrum at the end of acceleration is shown by the solid curve in Fig. 1(b). It can be observed that when the focal spot is 0.8 m, the cutoff energy of protons increases by 83.5% compared to the case with a 3 m spot, reaching 444 MeV. Although consuming a higher laser energy, the proton energy only reaches 242 MeV in the case with larger focal spot. This counterintuitive result indicates that reducing the laser spot size, even at a fixed intensity and lower total laser energy, enhances proton acceleration significantly. To elucidate the underlying physics, we analyze the electron acceleration dynamics for the 0.8 m case in Fig. 1(c) and the 3 m case in Fig. 1(d). When the laser interacts with the NCD plasma, the ponderomotive force expels electrons forward and laterally, forming a density spike at the laser front, which can be seen in the axial electron density distribution near x = 1 m in Figs. 1 (c) and 1 (d). In the case with a 0.8 m focal spot, a higher peak electron density can be observed, indicating stronger laser ponderomotive pushing on electrons. This results in more significant electron pile-up and more intense electron heating, as further evidenced by the electron energy spectrum represented by the dashed curve in Fig. 1(b).
@@ II.2 Down-ramp density target
shown as the red curve, where . For fixed areal density and laser parameters, protons with cutoff energies exceeding 700 MeV, a 60% increase over conventional targets, are generated with this down-ramp behind a 1.5-m-thick target, as evidenced by the spectrum in Fig. 4(b).

### v2

@@ Abstract
Enhancing proton energy is of great importance in laser-driven proton acceleration with finite laser energy for applications such as cancer therapy. We demonstrate an unusual acceleration scheme that achieves higher proton energies at lower laser energies by reducing the focal spot size. Through particle-in-cell simulations and theoretical modeling, we find that at small spot sizes (0.8 m), the proton energy is enhanced by 47.8%, much higher than that under conventional spot sizes (3 m). This is because the proton acceleration is dominated by electrons driven by an enhanced ponderomotive force at small spot sizes, generating stronger charge-separation fields that propagate faster. To further improve the proton energy, we analytically derive an optimal electron density profile, which enables phase-stable proton acceleration with an energy increased by 88.7%. These results are robust across parameter variations, suggesting that advanced focusing techniques and optimal plasma profiles could loose the requirement of laser energy, potentially reducing the dependence on large-scale laser facilities for medical and scientific applications.
@@ II Results
We perform a series of two-dimensional particle-in-cell (PIC) simulations using the EPOCH epoch code. A linearly polarized Gaussian laser pulse with a normalized amplitude of , a pulse duration of 30 fs and a wavelength of 800 nm is incident normally on a fully ionized hydrogen target with a density of 20 and a thickness of 7.5 m, where is the classical critical density. The simulation domain is a 40 m × 10 m rectangular box divided into 4000 × 1000 cells, with 20 macro-particles per cell per species. To investigate the effect of laser spot size, we vary the full width at half maximum (FWHM) of the laser focal spot from 0.8 m to 5 m, with representative cases of 0.8 m and 3 m selected for detailed analysis.
@@ II.1 Tightly focused laser
Figure 1(a) shows the temporal evolution of the maximum proton energy for the 0.8 m and 3 m spot sizes, and the proton energy spectrum at the end of acceleration is shown by the solid curve in Fig. 1(b). It can be observed that when the focal spot is 0.8 m, the cutoff energy of protons increases by 47.8% compared to the case with a 3 m spot, reaching 371 MeV. Although consuming a higher laser energy, the proton energy only reaches 242 MeV in the case with larger focal spot. This counterintuitive result indicates that reducing the laser spot size, even at a fixed intensity and lower total laser energy, enhances proton acceleration significantly. To elucidate the underlying physics, we analyze the electron acceleration dynamics for the 0.8 m case in Fig. 1(c) and the 3 m case in Fig. 1(d). When the laser interacts with the NCD plasma, the ponderomotive force expels electrons forward and laterally, forming a density spike at the laser front, which can be seen in the axial electron density distribution near x = 1 m in Figs. 1 (c) and 1 (d). In the case with a 0.8 m focal spot, a higher peak electron density can be observed, indicating stronger laser ponderomotive pushing on electrons. This results in more significant electron pile-up and more intense electron heating, as further evidenced by the electron energy spectrum represented by the dashed curve in Fig. 1(b).
@@ II.2 Down-ramp density target
shown as the red curve, where . For fixed areal density and laser parameters, protons with cutoff energies exceeding 700 MeV, a 88.7% increase over conventional targets, are generated with this down-ramp behind a 1.5-m-thick target, as evidenced by the spectrum in Fig. 4(b).